OpenStudy (anonymous):

f(x)+f(x+2)=root3 f(x+1) f(5)=3 so find f(5+12r) where r =0,1,2.........99

OpenStudy (shubhamsrg):

sqrt is only on 3 right ? not f(x+1)

OpenStudy (anonymous):

i think yes\[f(x)+f(x+2)=\sqrt{3} f(x+1)\]

OpenStudy (anonymous):

Dear Sir, Do you require a one-off special case solution or a METHOD to SOLVE THIS and MANY OTHER SIMILAR PROBLEMS ?

OpenStudy (anonymous):

The METHOD is well known its:

OpenStudy (anonymous):

u want to say its periodic?\[f(5+12r)=f(5)=3 \ \ \ r=0,1,2,...\]

OpenStudy (anonymous):

i've seen it before but i cant remember how to do it :(

OpenStudy (anonymous):

FOURIER TRANSFORM ( or a related technique of Laplace transform). AS is well known shift OPERATOR IN TIME DOMAUN is equivalent to PHASE MULTIPLIER IN FREAUENCY DOMAIN

OpenStudy (anonymous):

So let us begin the solution

OpenStudy (anonymous):

Let us TRANSFORM BOTH SIDES OF THIS EQUATION INTO THE FREQUENCY DOMAIN

OpenStudy (asnaseer):

isn't this a linear recurrence relation?

OpenStudy (asnaseer):

where you use a characteristic equation to solve it?

OpenStudy (anonymous):

Shifting by D like f(x+D) is equivalent to MULTIPLICATION by exp(-iDf)

OpenStudy (anonymous):

Of COURSE IN THE FOURIER DOMAIN

OpenStudy (anonymous):

SO your equation becomes:

OpenStudy (anonymous):

yes asnaseer its similar but there is only one given condition

OpenStudy (asnaseer):

I thought you only needed one given condition to solve a linear recurrence relation?

OpenStudy (anonymous):

F(f) + exp(-iDf)*F(f) = sqrt5 * exp(-if)F(f)

OpenStudy (anonymous):

as one can see this is a trivial linear equation on F(f)

OpenStudy (asnaseer):

for what its worth, this technique is described here: http://hcmop.wordpress.com/2012/04/20/using-characteristic-equation-to-solve-general-linear-recurrence-relations/

OpenStudy (anonymous):

by the way the generating function approach is equivalent - via Z-transform, which is a complex variant of Fourier transform

OpenStudy (anonymous):

F(f) (1 + exp(-i5f)) = sqrt(5)*exp(-if)

OpenStudy (anonymous):

F(f) =\[F(f) = \frac{ \sqrt{5}e^{-if}}{ 1 + e^{-i5f} }\]

OpenStudy (anonymous):

Now @gauti1989 this is a simple fraction - ITS INVERSE FOURIER TRANSFORM IS WELL KNOWN, and you can find it in any table and possibly even Wolfram Alpha will compute the inverse transform of this. Then you require that you condition will also be true and obtain the final answer

OpenStudy (anonymous):

If you want 'just the answer" without understanding the function - then notice that you have a simple recursion equation with 2 previous members of the sequence f(x+2)=root3 f(x+1) - f(x) which can be rewritten thus

OpenStudy (anonymous):

f(x+1) = root3 f(x) - f(x-1)

OpenStudy (anonymous):

So just jump forward 1 index, 2 indexes, 3 indexes - you will quickly discover a regularity

OpenStudy (anonymous):

Or , you can look up "z-transform" + "recursive equation" in ggl

OpenStudy (anonymous):

recursive equation will not work since it requiers 2 given condition

OpenStudy (anonymous):

i m not getting it xactly plz help me

OpenStudy (anonymous):

\[f(x)+f(x+2)=\sqrt{3} \ f(x+1)\]let x=x+1\[f(x+1)+f(x+3)=\sqrt{3} \ f(x+2) \ \ (1)\]let x=x-1\[f(x-1)+f(x+1)=\sqrt{3} \ f(x) \ \ (2)\] add (1) and (2)\[f(x-1)+f(x+3)+2f(x+1)=\sqrt3 \ (f(x)+f(x+2))=\sqrt3 \ (\sqrt{3} \ f(x+1))\]\[=3f(x+1)\]\[f(x-1)+f(x+3)=f(x+1)\]check till here

OpenStudy (anonymous):

\[f(x-1)+f(x+3)=f(x+1) \ \ (1)\]let x=x+2\[f(x+1)+f(x+5)=f(x+3) \ \ (2)\] add (1) and (2)\[f(x-1)+f(x+3)+f(x+1)+f(x+5)=f(x+1)+f(x+3)\]\[f(x-1)+f(x+5)=0\]

OpenStudy (anonymous):

\[f(x-1)+f(x+5)=0\]let x=x+1\[f(x+6)+f(x)=0\]

OpenStudy (anonymous):

\[f(x+6)=-f(x)\]let x=x+6\[f(x+12)=-f(x+6)=-(-f(x))=f(x)\]\[f(x+12)=f(x)\]so f is a periodic function with \(T=12\)

OpenStudy (anonymous):

\[f(5+12r)=f(5)=3 \ \ \ r=0,1,2,...\]

OpenStudy (anonymous):

i hope its clear :)